Find the sum: (-3/-11) + (5/9) Find the sum: (-3/-11) + (5/9)

Find the sum:    (-3/-11) + (5/9)
Find the sum:    (-3/-11) + (5/9)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= 3/11 + 5/9
Take the LCM of the denominators of the given rational numbers.
LCM of 11 and 9 is 99
Express each of the given rational numbers with the above LCM as the common denominator.
Now,
(3/11)= [(3×9)/ (11×9)] = (27/99)
(5/9)= [(5×11)/ (9×11)] = (55/99)
Then,
= (27/99) + (55/99) ... [∵denominator is same in both the rational numbers]
= (27 + 55)/99
= (82/99)

Thanks